# High-Fidelity and Complex Test Data Generation for Google SQL Code Generation Services

### v3

@@ I Introduction
The generation also needs to ensure data value diversity to effectively simulate the wide range of scenarios encountered in real-world systems, including both common use cases and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [17, 16].
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as database schema and protocol buffers. For clarity, throughout this paper, we use the term ”schema” to generally refer to a formal structural definition, encompassing both Google SQL schema definitions and/or Protocol Buffer (proto) definitions with potentially nested fields. LLMs can generate mock data that more closely approximates real-world data distributions and patterns based on their acquired real-world knowledge. LLMs can capture subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [18, 19]. This is crucial for comprehensive production testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [20, 21, 19, 22].
@@ II SQL code generation services
In this section, we examine the challenges of generating high-fidelity test data for two real-world GoogleSQL [13] SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
@@
| Generation time per instance (50p) | 1 hour222An engineer, familiar with the table schema, was tasked with generating a test data instance within one hour. While individual speed may vary, it is practically impossible to curate a comprehensive instance in such a limited timeframe. | 37 sec | 167 sec |
@@ II-B NL2SQL
II-C Problem Definition
@@ III-A Pre-processor
Handling Subqueries:
The pre-processor module treats each subquery as a separate query and associates each of the signals for data generation mentioned above separately for each subquery and the main query, as illustrated in Figure 4.
@@ III-B Context extractor
This module generates context relevant to the schema. We use a separate system to collect human written comments for a schema and proto from comments in data schema, source code and documentations. These comments are available at the schema/proto level and also individual column level. We use LLM to make these comments consistent and fill in missing columns.
For protos, by themselves or as a column in a schema, we generate the proto description with comments for additional context for data generation.
This component will also be able to extract column correlations from the SQL query body and annotated schema that would be useful in generating high-fidelity data.
Minimizing edge cases with constraints:
We pass constraints extracted directly from the test SQL query, as illustrated in Figure 4, so that the model can explore and generate different values a relevant column in the predicates should take for the test (for instance, the LLM understands that it should generate for currency from the query in Figure 1). The same set of constraints are later used for the semantic quality verification of the generated data as well. In Section VI-B we discuss another example where our LLM-based generation successfully generate test data around constraint, leaving no quarter without test data.
@@ III-D Post-processor
This module can add the data constraints and variations by copying the basic data generated by the LLM or existing data written by the user. The extracted join constraints in the pre-pro in the pre-processing are also enforced by the post-processor, where the column values in the joined tables at the join columns are matched and copied from one table data to another. Since the constraints and variations will affect only a smaller number of columns in a proto, we could easily just modify those components.
This module will not make any LLM calls and will update the relevant fields identified in the preprocessor module. This will also provide us a chance to correct any problems with the generated data. This will also reduce the total time taken for data generation as the LLM call latency scales linearly with the number of data points to be generated. This will also make the changes more deterministic in nature.
@@ III-E Data validator
We would evaluate the generated data using a proto parser [10]. This has an in-built check for the datatype and any other constraint on the proto fields. In case of any failure, we would be retrying the data generation, at most 2 retires, along with an error message. The integrated testing will be done by the framework.
@@
Semantic constraints and validation is not easy to capture with conventional approaches, and we use the LLM for semantic validation (LLM-as-a-judge [31]) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 8. Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks (Table VI).
@@ IV-A Evaluation setup
For our test data generation pipeline, we used the latest gemini-flash-2.0 model and with 10 concurrent threads handling nested structures (e.g., nested field that leads to another schema/proto structure). The threading happens at the same level of nesting to ensure that there is no data dependency between concurrent threads.
@@ IV-B Test coverage with more complex SQL query testing
Table V presents the ablation study on NL2SQL test data generation by removing each component from our data generation pipeline. Column grouping and data validation steps are more relevant for semantic validity and the test coverage by test executability remains unchanged without them. We do not include SQL Code Assistant, since the unit test cases consists of simple selection and projection queries – SQL parsing and schema fetching are sufficient to populate the relevant schema structures and fields to get the full coverage improvement.
@@
| Test coverage (%) | +12.58 | +69.59 |
| Test accuracy (%) | +0.0 | +20.4 |
@@
| Post-processing | |
| Data Validation |
@@ IV-C Generation quality evaluation with complex nested data structures
Next, we evaluate the completeness and correctness of the generation for complex nested data formats. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types. Figure 9 shows the normalized results (0 to 1, higher is better), using key criteria from Table II – we excluded the criteria for primary key generation, which is always true for our test data. Despite issues with some inaccessible schemas in production environment, our data generation successfully produces test data compliant with complex, real-world nested structures.
@@ V Related works
Despite emerging interest in using LLMs for test data generation [14, 23, 24, 25], existing approaches often fail to meet the complex requirements of industrial testing. While LLMs can handle basic schema adherence [21], out-of-the-box models struggle with subtle data relationships [18] and become incomplete for large, nested schemas (Table II).
@@
To address these limitations, we propose a complex test data generation method based on the pre-trained LLM, gemini, requiring no post-training.
@@ VI-A Scaling and cost of the generation
While much more efficient and scalable than manual test data generation, as illustrated in Table II, our test generation process can easily be deployed and parallelized across different batch of test queries/schemas, and also across different nested structures/fields within the same schema. The test data generation efficiency is important, also the main pain-point shared by the practitioners (e.g., developers submitting test queries and manual test data), as the data generation time takes much longer than the actual test generation/execution time.
@@
| Input tokens | 162199.5 ( 350852.1) | 57082.3 ( 102381.8) |
| Output tokens | 3890 ( 3121.5) | 8054.2 ( 13040.1) |
| Generation time (sec) | 66.0 ( 53.6) | 57.5 ( 64.36) |
Table VII presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model [36] (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
@@ VI-B Generating for test SQL query semantics
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query [37, 38]. While the NL2SQL prediction quality evaluation is beyond the scope of this study, high-fidelity test data should not only be adhere to the underlying schema, but also semantically relevant to the target SQL query (natural language question). One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section IV-E.
To illustrate the problem, consider a business analytic query in Figure 11. High-fidelity test data for this query should ensure that enough example conversions instances are present across all the quarters in the requested date range. Our LLM-based generator can recognize such semantic importance of date_trunc(logdate, quater) and generate data for all quarters in the range. Moreover, if one has a domain knowledge about the data, such as “conversions were better in Q4 than Q1,” then it can be passed to the generator as user_criteria as part of input, which can guide the generation to favor Q4 instances. Such understanding and reasoning capabilities of LLMs can scale semantic-aware test data generation.
@@ VI-C Generation error analysis
We observe some common error patterns in our data generation process:
Output data formatting error. We request output in JSON format, but the actual responses had incorrect JSON formatting failing the JSON parser. This can be fixed with retries.
Missing fields from large and complex data format, which requires some post-processing steps to identify and retry.
Generating fewer than requested amount of data/rows, which can be fixed with retries to generate data incrementally.
For some schemas with a very small number of primitive columns, the LLM hallucinated and introduced new columns. Care is needed to keep track of requested columns/fields for each (nested) schema.
Missing some edge cases relevant to given SQL queries/functions.
We have implemented some number of safeguards against LLM hallucination. Most notably, we have observed in some cases with simpler data formats, the LLM introduced new fields – even nested ones. Figure 12 illustrates one such example where many more new fields are introduced next to just 3 required fields; the LLM even introduces nested fields.
Addressing the issue of missing edge cases relevant to specific SQL queries is challenging, as we lack an oracle (or access to production data). While our pipeline uses an ”LLM-as-a-judge” for semantic validation, this risks false positives. To improve fidelity, we could integrate domain knowledge (as discussed in Section VI-B) to guide the LLM toward specific cases. This remains an important problem for improving mock data generation in production. Furthermore, techniques for statistically imputing missing values and estimating their impact on query results [39, 40] warrant future investigation in this setting.
@@
In this paper, we addressed the critical challenge of generating semantically relevant and complex test data for Google SQL code generation services, particularly where production data access is restricted. We highlighted the limitations of traditional data generation methods that were used by developers, focusing on their inability to handle large, deeply nested data structures, like protocol buffers, and to maintain semantic integrity of the target test cases. By leveraging the latest Gemini 2.5 Flash model with strategic pre- and post-processing, we successfully generated comprehensive test data, compliant with large, complex data structures. This data effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries, and significantly improves test coverage. Our experiments demonstrate a reliable and practical method for generating high-fidelity test data, which is essential for industrial environments where access to production data is restricted. While our current implementation targets Google SQL, the underlying methodology is dialect-agnostic.

### v4

@@ I Introduction
The generation also needs to ensure data value diversity to effectively simulate the wide range of scenarios encountered in real-world systems, including both common use cases and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [28, 16, 33, 29].
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as database schema and protocol buffers. For clarity, throughout this paper, we use the term ”schema” to generally refer to a formal structural definition, encompassing both Google SQL schema definitions and/or Protocol Buffer (proto) definitions with potentially nested fields. LLMs can generate mock data that more closely approximates real-world data distributions and patterns based on their acquired real-world knowledge. LLMs can capture subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [42, 39, 4]. This is crucial for comprehensive production testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [36, 4, 39, 31].
@@ II SQL code generation services
In this section, we examine the challenges of generating high-fidelity test data for two real-world GoogleSQL [14] SQL code generation services: SQL Code Assistant and NL2SQL. A critical requirement for testing these services is the ability to handle the rigid structural constraints of GoogleSQL, which supports deeply nested types like array, struct, and proto. Traditional constraint-based generators [28, 16, 33, 29], often built for flat, relational schemas, cannot natively parse these complex nested structures. Because these tools lack the logic to maintain semantic integrity across multiple levels of nesting, they are unable to produce the syntactically valid instances required for executing the SQL functions under test.
@@
| Generation time per instance (50p) | 1 hour222A domain-expert engineer required over one hour to manually curate a single, comprehensive test instance. Given the schema complexity, manual generation is practically infeasible for large-scale industrial testing. | 37 sec | 167 sec |
@@ II-B NL2SQL
(a) Natural Language Question
(b) Golden GoogleSQL Query
@@
II-C Testing Requirements and Evaluation Methodology
Industrial SQL code generation services require more than just syntactically valid test data; they demand a high degree of semantic fidelity to ensure evaluation validity. The primary evaluation metric for these systems is Execution Accuracy (EX), which assesses the model’s performance by executing both the predicted SQL and the golden SQL on a database and comparing their output result sets [7, 11]. For this evaluation to be meaningful, the underlying test data (generator) must satisfy the following critical requirements. First, due to strict security policies, the generator cannot have access to production data samples. Instead, it must build the right context by synthesizing metadata, including source code comments, schema definitions, documentation, and the test SQL queries/questions themselves. Second, the data must be generated with values that satisfy the predicates in the SQL queries, so that the result sets are not empty. Semantically irrelevant data can lead to empty result sets for both predicted and golden queries, producing trivial matches that undermine the evaluation’s validity.
II-D Problem Definition
@@ III-A Pre-processor
(a) Input SQL Query
(b) Extracted Generation Targets
(c) Extracted Join Constraints
Handling Subqueries and Table-Level Parallelism:
Our framework decomposes complex SQL queries into atomic data generation targets based on the underlying table schemas. This allows for the parallelization of the generation process across independent tables referenced in the main query and its nested subqueries.
To maintain semantic integrity across the entire query tree, the pre-processor extracts data generation signals (e.g., predicate constraints) and maps them to each referenced table. Independent tables are flagged as primary generation targets under “Subqueries:” in the pre-processor output, while redundant or dependent references are deduplicated to optimize token usage and generation time (see Figure 4 for an illustration). For multi-table queries involving joins, the pre-processor isolates join-key associations, which are subsequently enforced during the deterministic post-processing phase to ensure relational consistency.
@@ III-B Context extractor
The context extractor module generates context relevant to the schema. Industrial schemas often suffer from cryptic (non-human readable) identifiers and a lack of formal CHECK constraints, creating a significant challenge for traditional rule-based generators. To bridge this gap, we use a separate system to aggregate human-written metadata, including schema descriptions, source code comments, and technical documentation. These comments are available at the schema/proto level and also individual column level. Because this information is often fragmented and under-specified, we use the LLM to make these comments consistent and fill in missing column descriptions [30, 21, 23, 34]. This interpolation task is essential for our zero-shot setting (i.e., no production data seeding to learn the distributions). While we acknowledge that LLM-generated annotations are not perfect, they provide the necessary semantic grounding to generate executable and coherent test data where other tools fail.
Beyond metadata, the module extracts functional signals directly from the test SQL query, as illustrated in Figure 4, so that the model can explore and generate different values a relevant column in the predicates should take for the test (for instance, the LLM understands that it should generate for currency from the query in Figure 1). The same set of constraints are later used for the semantic quality verification of the generated data as well. In Section VI-B we discuss another example where our LLM-based generation successfully generate test data around constraint, leaving no quarter without test data.
@@ III-C Prompt builder
(a) Without Column Grouping
(b) With Column Grouping
(a) Source Context (Human Comments & Docs)
(b) Without Annotation Generic/Hallucinated
(c) With Annotation Semantically Correct
@@ III-D Post-processor
This module functions as a deterministic layer to enforce data semantic and relational constraints on the raw data generated by the LLM. It specifically targets two areas:
Predicate and Data Constraint Enforcement: The module applies user-defined signals and extracted predicate constraints (e.g., specific date ranges or categorical filters). Since these constraints typically affect only a small subset of the schema, the post-processor performs direct value substitution or correction on the generated instances.
Relational Join Integrity: Using the associations (join constraints) extracted by the pre-processor, this module synchronizes join-key values across tables. Values from a primary table are matched and copied to the corresponding columns in joined tables, ensuring that the generated mock data remains join-consistent even when the LLM’s initial output lacks relational coherence.
By decoupling these deterministic operations from the probabilistic generation of the LLM, we significantly reduce total latency—as these updates do not require additional LLM calls – and provide a robust mechanism to correct structural or semantic inconsistencies.
It is also important to note that enforcing join constraints during post-processing allows us to decouple the generation process at the table or schema level. By treating each table as an independent generation target, we can maximize horizontal parallelization across the workload.
@@ III-E Data validator
Our framework employs a dual-stage validation pipeline to ensure both structural integrity and semantic coherence. We validate the syntax of the generated data instances using a static parser, like proto-parser [9]. This deterministic parser ensures that the generated data adheres to the schema and satisfy all the type and structural constraints.
Because conventional static parsers cannot capture complex business logic or test SQL-specific predicates for semantic coherence (to the test question and to other column values), we also utilize an LLM-as-a-judge [44] approach. The validator receives the generated data alongside the semantic constraints extracted by the pre-processor (e.g., data ranges or current types). To ground the model’s reasoning and resolve inconsistencies, we implement a retry mechanism. If a generation fails at validation stage, the system retries up to three times, feeding the specific error message or semantic failure reasons back into the generator to guide the next attempt. This iterative grounding reduces semantic error rates (Section IV-D). The template in Figure 8 illustrates the detailed instructions for the judge.
@@
Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks (Table VI).
@@ IV Evaluation results
Schema Format Versatility: We evaluate the methodology’s robustness by applying it to JSON, a format with deeply nested structures comparable to GoogleSQL’s nested types, confirming the approach is not limited to GoogleSQL and Protocol Buffer definitions
@@ IV-A Evaluation setup
While we provide a preliminary comparison in Table II, our primary evaluation focuses on the effectiveness of our proposed pipeline. We found that “Manual” curation is practically impossible for large-scale industrial testing due to the time required (over an hour per test data instance), and “Vanilla LLM” fails to generate syntactically valid data for deeply nested structures. Furthermore, existing statistical and LLM-based tabular generators [42, 4] were excluded as they require production data samples—which are strictly restricted in our environment—to learn the data distributions. Therefore, our method serves as a necessary zero-shot solution where prior tools fail to function.
For our test data generation pipeline, we used Gemini 2.5-Flash and with 10 concurrent threads handling nested structures (e.g., nested field that leads to another schema/proto structure). The threading happens at the same level of nesting to ensure that there is no data dependency between concurrent threads.
@@ IV-B Test coverage with more complex SQL query testing
Table V presents an ablation study for the NL2SQL pipeline, quantifying the relative impact of each component on test coverage. Removing the Pre-processor or Schema Fetching results in a 100% loss in coverage, as these modules provide the foundational metadata required for any data generation. The removal of Schema Annotations leads to a 47.76% decrease in coverage, highlighting the importance of user comments and documentation-derived context for complex, real-world schemas. While Column Grouping and Data Validation show a 0% impact on raw test coverage, these components are critical for maintaining the semantic validity and logic required for representative evaluation results; without them, test executability remains unchanged, but data integrity is compromised. Removing the Post-processor reduces coverage by 20.15%, demonstrating the necessity of deterministic join enforcement for queries involving multi-table dependencies.
We omit SQL Code Assistant from this ablation because its test cases primarily involve simple selection and projection queries; for these scenarios, basic SQL parsing and schema fetching are sufficient to achieve full coverage improvement. It is important to note that this is still infeasible with manual curation or existing static data generation tools. The LLM-based approach is essential because traditional tools cannot reason through the extreme structural complexity and deep nesting inherent in our production schemas.
@@
| Coverage (%) | +12.58 | +69.59 |
| Accuracy (%) | +0.0 | +20.4 |
@@
| Post-Processor | |
| Data Validator |
@@ IV-C Generation quality evaluation with complex nested data structures
Next, we evaluate the completeness and correctness of the generation for complex nested data formats. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types. Figure 9 shows the normalized results (0 to 1, higher is better), using key criteria from Table II – we excluded the criteria for primary key generation, which is always true for our test data. Despite issues with some inaccessible (or stale) schemas in production environment, our data generation successfully produces test data compliant with complex, real-world nested structures.
@@ IV-D Semantic data constraints validation with LLM-as-a-judge
IV-F Generalizability to Other Data Formats
| Model | Format | Detailed | Strict |
| Our Method (Gemini 2.5-Flash) | 100.0 | 92.11 | 59.86 |
| Claude Sonnet 4 | 99.05 | 90.73 | 57.90 |
| Magistral Medium 2506 | 98.10 | 90.36 | 59.81 |
| DeepSeek R1 0528 | 97.90 | 89.57 | 57.33 |
| Gemini 2.5 Pro | 97.52 | 89.00 | 56.19 |
| Qwen3 235B A22B | 97.14 | 88.33 | 56.19 |
To demonstrate the generalizability of our approach beyond GoogleSQL and proto, we evaluated our pipeline using the DeepJSONEval benchmark [45]. This dataset focuses on generating deeply nested JSON objects that must strictly adhere to complex schemas and semantic constraints. As shown in Table VII, our method leveraging Gemini 2.5-Flash achieves the best results, even outperforming those by more advanced SOTA models. This result confirms that our pre- and post-processing modules effectively handle diverse nested structured data formats without dialect-specific tuning.
@@ V Related works
Despite emerging interest in using LLMs for test data generation [15, 27, 20, 26], existing approaches often fail to meet the complex requirements of industrial testing. While LLMs can handle basic schema adherence [4, 30, 21, 34, 19, 23], out-of-the-box models struggle with subtle data relationships [42] and become incomplete for large, nested schemas (Table II).
One of the major barriers to using existing LLM-based tabular generators in our setting is their reliance on actual table samples to learn data distributions (e.g., fine-tuning). Our approach addresses this gap by operating in a true cold-start regime. Instead of requiring seed data, we leverage the LLM’s internal reasoning by using the natural language question and the SQL query body as primary signals to generate semantically relevant data.
@@
To address these limitations, we propose a complex test data generation method based on the pre-trained LLM, Gemini, requiring the testing function/SQL statements as well as the corresponding business logic questions (NL2SQL).
@@ VI-A Scaling and cost of the generation
| Input tokens | 162,199.5 ( 350,852.1) | 57,082.3 ( 102,381.8) |
| Output tokens | 3,890.0 ( 3,121.5) | 8,054.2 ( 13,040.1) |
| Time (sec) | 66.0 ( 53.6) | 57.5 ( 64.36) |
| Metric | Temperature | |||
| 0.1 | 0.5 | 1.0 | 1.8 | |
| Unique instances at 10k | 10,000 | 10,000 | 10,000 | 10,000 |
| video_sid uniqueness | 50% | 50% | 40% | 70% |
| country_code coverage | 87.14% | 7.2% | 39.75% | 44.17% |
| Enum field coverage | 100% | 100% | 100% | 100% |
Our automated pipeline is highly parallelizable across query batches and deeply nested structures, addressing the primary industrial bottleneck where manual data curation time often exceeds test execution time (Table II). Table VIII quantifies the efficiency leap: while a domain expert requires over an hour to curate a single comprehensive instance for complex schemas, our framework generates high-fidelity data in 57.5–66.0 seconds on average. The high variability in input tokens reflects the diverse schema complexities, including recursive nested structures, successfully handled by the system without production data access.
To assess scalability for large-scale workloads, we evaluated the system’s ability to maintain diversity across 10,000 generated instances (Table IX). The pipeline sustains high quality without significant repetition as dataset size increases. For high-cardinality fields like video_sid, the model achieved up to 70% uniqueness. While most SQL validation tasks require only a few hundred unique identifiers, our framework supports deterministic post-processing—such as ID tracking or incremental assignment—to ensure absolute uniqueness for datasets exceeding 10,000 rows. For constrained fields like country codes, the LLM explored up to 87.14% of the value space, while enum coverage remained perfect (100%) across all temperatures. These results confirm that the method produces semantically grounded, diverse datasets suitable for industrial-scale SQL testing.
@@ VI-B Generating for test SQL query semantics
Semantic coherence between test data and the target SQL query is essential for evaluating SQL code generation services like NL2SQL. High-fidelity test data must move beyond structural integrity to reflect the functional logic of the query. For instance, the business analytics query in Figure 11, which calculates quarterly conversion rates, requires data instances distributed across every quarter of 2022 to produce a valid report. Our LLM-based generator can recognize such semantic importance of date_trunc(logdate, quater) and generate data for all quarters in the range. Moreover, if one has a domain knowledge about the data, such as “conversions were better in Q4 than Q1,” then it can be passed to the generator as user_criteria as part of input, which can guide the generation to favor Q4 instances. Such understanding and reasoning capabilities of LLMs can scale semantic-aware test data generation.
@@ VI-C Generation error analysis
Our evaluation revealed several recurring error patterns in the generation process: formatting inconsistencies (invalid JSON), structural omissions (missing fields or rows), and hallucinations where extraneous columns were introduced—particularly in simpler schemas. Most of these issues are effectively mitigated through automated retries and schema-tracking safeguards that enforce strict adherence to the requested field list.
Addressing missing edge cases for complex SQL predicates remains a challenge without production data or an oracle. While our LLM-as-a-judge validator provides semantic verification, it is susceptible to false positives. To improve fidelity, we recommend integrating domain-specific constraints (as discussed in Section VI-B) to guide the LLM toward relevant edge cases. Future investigation into statistical imputation and estimation techniques [3, 8] could further enhance mock data quality in production environments.
@@
In this paper, we addressed the critical challenge of generating semantically relevant and complex test data for Google SQL code generation services, particularly where production data access is restricted. We highlighted the limitations of traditional data generation methods that were used by developers, focusing on their inability to handle large, deeply nested data structures, like protocol buffers, and to maintain semantic integrity of the target test cases. By leveraging the Gemini 2.5-Flash model with strategic pre- and post-processing, we successfully generated comprehensive test data, compliant with large, complex data structures. This data effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries, and significantly improves test coverage. Our experiments demonstrate a reliable and practical method for generating high-fidelity test data, which is essential for industrial environments where access to production data is restricted. While our current implementation targets Google SQL, the underlying methodology is dialect-agnostic.
Acknowledgment
We would like to thank John Cieslewicz, Kashmira Phalak, Yannis Papakonstantinou and the anonymous reviewers for their valuable feedback. We also thank Yuchen Zhang for his contribution in the early stages of this project.